1. Paula is 6 years old. Her uncle is 4 times as old. How old will Paula be when she is half as old as her uncle?
Please explain and answer.. thanks

Answer by checkley71(8403)

P=6 U=4*6 OR U=24 NOW
IN X YEARS P WILL BE 6+X AND HER U WILL BE 24+X AND P WILL BE 1/2 U
OR (6+X)=1/2(24+X) OR 6+X=12+.5X OR .5X=6 OR X=6/.5 OR X=12 THUS IN 12 YEARS P WILL BE 6+12 OR 18 AND HER U WILL BE 24+12 OR 36 AND 18=(1/2)*36 OR 18=18